OpenStudy (anonymous):

Help with finding the derivative of y=x(6^-2x) and how to work it out?

OpenStudy (anonymous):

\[y=x(6^{-2x})\]

OpenStudy (anonymous):

use uv rule

OpenStudy (anonymous):

d(x(6^-2x))/dx=x*d/dx(6^-2x)+6^-2x d/dx(6^-2x)=(6^-2x)*-2*log6

OpenStudy (anonymous):

could you use the product rule?

OpenStudy (anonymous):

yes you can uv rule is the same as product rule

OpenStudy (anonymous):

oh okay